# Privacy Vulnerabilities in Marginals-based Synthetic Data

### v1

@@ I Introduction
SDG techniques with built-in privacy assurances have however been shown to leak private information from the training data [5, 12, 18]. One fundamental way of showing leakages is to conduct a membership inference attack (MIA) on the synthetic data, the goal being to glean which records were present in the hidden training data [16, 35]. Practically speaking, a successful MIA on cancer-related synthetic data or AI model could determine an individual’s medical history, for example.
@@ II-A Synthetic data generation
Many types of SDG algorithms have been developed to generate data in different ways. One prominent class of SDG employs generative adversarial networks (GANs). Algorithms using this paradigm perform quite well when generating images (e.g. [34, 37]). However, they tend not to produce high quality tabular data [27, 32]. The other main approach to SDG is to instead measure statistical properties from the training data $D_{train}$ and preserve them in the synthetic data $D_{synth}$. Specifically, we focus on “marginals-based” SDG algorithms, i.e. those that maintain marginal probability distributions to keep this likeness. [^5] It is worth briefly mentioning that similar statistics-preserving SDG algorithms rely indirectly on marginals, via copulas to capture the dependencies between features. But we don’t study copulas algorithms here. Additionally, some algorithms might be classified as projection-based [17]. However, many of these still operate by measuring marginals [2, 36] and using projection to preserve them. When it comes to tabular data, marginals-based SDG techniques are dominant in achieving the highest utility for machine learning applications [27, 32]. Naturally, a higher level of scrutiny must accompany their acclaim.
In this paper, we use our novel MAMA-MIA attack to investigate vulnerabilities of three state-of-the-art SDG algorithms that employ marginals in some way: MST [24], PrivBayes [40], Private-GSD [21] (and we include an investigation on RAP [2] in the appendix). What is notable about these algorithms, unlike several popular others, is that they each provide robust differential privacy (DP) [8] guarantees to protect the training data. So to maximize the inherently-conflicting aims of achieving high quality and strong privacy, these algorithms use marginals and apply DP, respectively. Descriptions of MST, PrivBayes, RAP, and Private-GSD are included in Appendix  III-C.
@@
$\hskip-8.0pt\hat{P}_{D}(v_{c}\mid\mathbf{v})=\frac{|\{d\,|\,d\in D_{\mathcal{P}}\wedge V_{c}(d)=v_{c}\ \wedge\mathbf{V}(d)=\mathbf{v}\}|}{|\{d\,|\,d\in D_{\mathcal{P}}\wedge\mathbf{V}(d)=\mathbf{v}\}|}$ (2)
@@
Differential privacy (DP) The well-known definition of DP states that a randomized algorithm $\mathcal{A}$ with range $O$ is $\varepsilon$-DP if for any two adjacent datasets $D_{1}$ and $D_{2}$ (datasets that differ in one entry)
@@ II-B Membership inference
Threat model Each MIA makes assumptions about what an adversary has access to, known as the “threat model”. Some MIAs use a threat model where an adversary knows little. Others grant a generous amount of knowledge to an adversary––the motivation being to drive the development of stronger privacy protections against stronger opponents. Our attack follows the latter inspiration. In our threat model, we use the “the auxiliary data assumption”, which admits knowledge of the population data to the adversary [16]. This is a generally accepted assumption [35, 1], though is argued against in [11].
@@
An even stronger threat model could grant “white-box” knowledge, where an adversary knows the internal weights and randomness applied by the generator. But we do not use this liberal assumption in our attack. There also is black-box access to the generator, where an adversary can execute the generator to create synthetic datasets without knowledge of what algorithm was used (a slightly weaker assumption than the one we use).
@@ II-C Related MIAs
References [4] and [31] describe attacks that rely on exhaustive “shadow-modelling” to learn about the hidden data. The idea behind these is to simulate the SDG algorithm on many random samples of the population data, thereby generating many “shadow datasets”, $\hat{D}_{synth}$. By generating many shadow synthetic datasets, it may be possible to notice differences in those that were trained using the target (i.e. the example the attacker is interested in) versus those that were not, informing whether or not the target was a member of the actual hidden training data $D_{train}$. Clearly, these attacks use both the “auxiliary data assumption” as well as the black-box access to the generator assumption (but not necessarily knowledge of the underlying SDG algorithm) to conduct the shadow modelling. And both works show superior inference accuracy using this method. However, it tends to require generating several hundred shadow datasets, using scads of compute resources.
@@
The TAPAS library [^6] https://tapas-privacy.readthedocs.io/en/latest/index.html provides a framework for several types of MIAs based on different threat models [16], including shadow modelling such as “Groundhog” attacks [31], distance-based attacks [22, 38], probability estimation attacks [19, 14], and others. In both their analysis and in [25], shadow modelling attacks using random queries yield the highest accuracy when using the auxiliary data and blackbox access to the generator assumptions. So in this paper, we provide a comparison of our attack with this leading MIA. (Meeus et al. find even greater success by smartly choosing which target records are most susceptible to attack [25]. We do not compare our work against this specialized approach, since ours is less restricted, and randomly samples which targets to attack.)
We also conduct a comparison of our attack against the DOMIAS MIA [35]. DOMIAS is highly relevant because, in addition to making clever use of the auxiliary data assumption (expounded upon in Section III), it outperforms several other MIAs [5, 12, 14]. But it does not use the black-box knowledge assumption, and is therefor generic to SDG algorithms.
@@ III MAMA-MIA: the MArginals Measurment Aggregation-based Membership Inference Attack
MAMA-MIA constructs a novel density estimation $\zeta$ in place of $\Lambda$. $\zeta$ incorporates the normalization idea from DOMIAS, but is tailored to which marginals-based algorithm was used, and therefor has a much stronger ability to detect characteristics of the training data. Applicable to any marginals-based SDG algorithm, the heuristic behind MAMA-MIA is to:
@@ III-C Adaptation of MAMA-MIA to each SDG algorithm
Like MST, it uses the conditionals-based graphical approximation in order to generate $D_{synth}$, [^8] This is done by sampling values directly from the conditionals, and then following the corresponding path along the directed graph. making this approximation highly relevant to an adversary. Edges that yield high mutual information are preferred by PrivBayes, but how many parents are allowed in each conditional depends on $\varepsilon$.
@@
A note on generalizability and efficiency Our density estimation approach can be tailored to similar SDG algorithms that estimate the probability distribution of a training dataset via marginals or other similar statistics, like conditionals or copulas. In Appendix section III-C, for example, we explain how MAMA-MIA can be adapted to RAP, another recent marginals-based SDG algorithm, and provide experimental results for it. [^9] However, we don’t include it in the main text because, as we demonstrate, we found it to produce data of poor quality, rendering any attack on it ineffective. Additionally, MAMA-MIA is highly efficient, and operates in linear time with respect to the size of $D_{synth}$ and $D_{aux}$ (we offer an asymptotic analysis in Appendix A-D).
@@ IV Experiments and results
We perform a series of experiments on MAMA-MIA to show its leading accuracy, and to demonstrate its far superior efficiency compared to state-of-the-art MIAs. These experiments include attacking each SDG algorithm – MST, PrivBayes, and Private-GSD – under various configurations.
@@ IV-A Experiment A results
Results are shown in Figure 3. Our MAMA-MIA attack’s accuracy is at par with or better than the other MIAs on data generated by all three SDG algorithms. Not surprisingly, the accuracy improves when the privacy budget $\epsilon$ is larger, i.e. when the privacy guarantees offered by the DP SDG algorithms are lower. DOMIAS did not perform significantly above random guessing, neither with BNAF nor with KDE, the two density estimators used by Van Breugel at al. [35]. [^13] There were settings in which DOMIAS performed better than random guessing, but still far worse than MAMA-MIA, and only when the datasets used were small (see appendix section A-B). We also provide a tabled version of our full results, with standard deviations, in the Appendix A-E.
But the striking advantage our MAMA-MIA technique has over its fellow MIAs is its efficiency. Shown in Table I are the average runtimes of each of the MIAs in experiment A. These times include the full attack pipeline on the set of targets; that is, both the shadow modelling stage and the inference stage of the MAMA-MIA attack, and the TAPAS attack. Recall from Section III-A, we only shadow model 50 times for MAMA-MIA, where we terminate the SDG process early (as soon as the focal-points are determined). This is particularly relevant for SDG algorithms that are computationally intensive, like Private-GSD, where the TAPAS shadow modelling step entails 500 complete executions. [^14] The times given are processor time, and so, with parallel processing, these times can be shortened. For DOMIAS, there is no shadow modelling. But even so, when using BNAF, DOMIAS takes the longest (except for on Private-GSD).
TABLE I: Total pipeline runtimes for one trial in experiment A, where $|D_{train}|=1,000$ records of SNAKE data.
@@ IV-D Additional results
We only conduct the TAPAS attack on Private-GSD for when $\varepsilon=10$ because of the high cost of generating 500 shadow datasets for each setting. [^16] Importantly, we reuse the 500 shadow datasets for all 30 trails under each setting. To give a sense, just generating the Private-GSD shadow datasets for one TAPAS attack took 387 days of processor time. The second part of the attack that processes the shadow datasets took 19 minutes.
@@
The non-monotonicity of the qualities of Private-GSD housing data in Figure 9 part (C) is somewhat puzzling, and could be the result of an incompatibility between what our distance measurement captures and what Private-GSD captures on the binned housing data. The far-off distances in Private-GSD’s data in part (b) of the figure, and the odd spike in its distance in part (d), show how the quality of Private-GSD data degrades when only using 20,000 generations (the default amount) when $|D_{train}|$ and $|D_{synth}|\geq 3,000$ records. However, when 100,000 generations are used, the quality is as good as data produced by PrivBayes. We always allow for early stopping of Private-GSD, which can explain why, when $|D_{train}|$ and $|D_{synth}|\leq 1,000$ records, using 20,000 generations versus 100,000 generations yields data with practically identical distances.
So because of this, in all experiments except for experiment C, we set Private-GSD to use 100,000 generations. [^18] Because of the computational load of Private-GSD, in experiment C, where we generate 4,000 shadow datasets for one TAPAS attack, we use its default 20,000 generations instead of 100,000. This appears to be completely adequate for experiment C. That is, when $|D_{train}|=1,000$, the respective dataset qualities will be nearly identical under either configuration (shown in Figure 9).
$\textit{MA}=(\textit{tpr}-\textit{fpr}+1)/2$ (4)
where tpr, fpr are the true positive rate and false positive rate, computed by weighting individual predictions by their distance from a $0.5$ threshold: $2\cdot|0.5-p_{i}|,p_{i}\in P$. Like with AUC, an MA score of $0.5$ is as good as random guessing. In all of our experiments, MA almost always ended up being within $\pm$ 0.02 of what the AUC score was. Our accuracies are noticably better here than for the above experiments because the objective of this challenge was “set MI” (where targets are grouped into sets of five or more records, and either the entire set of records is in $D_{member}$ or none are). We provide more set MI results in the appendix section A-E.
@@ A-C Experimental parameters
TABLE V: Average AUC scores ($\pm$ one standard deviation) over 30 trials for MAMA-MIA and TAPAS attacks on the SNAKE data. $D_{train}=1,000$, and TAPAS uses 500 shadow datasets.
@@
TABLE VI: Average AUC scores ($\pm$ one standard deviation) over 30 trials for MAMA-MIA and TAPAS attacks on the housing data. $D_{train}=1,000$, and TAPAS uses 500 shadow datasets.
@@
For MST and PrivBayes, we use the implementations given in the Reprosyn library. [^20] https://github.com/alan-turing-institute/reprosyn For Private-GSD and RAP, we use the authors’ implementations. [^21] https://github.com/giusevtr/private_gsd, [^22] https://github.com/amazon-science/relaxed-adaptive-projection
For all experiments (except in experiment C because of computational limitations), we increase the default number of generations Private-GSD uses from 20,000 with early stopping to 100,000 with early stopping. As shown in Figure 9, this only makes a difference on the data quality for when $D_{train}$ and $D_{synth}$ have more than 1,000 rows. Therefor, this parameter change did not impact the accuracy of MAMA-MIA in experiment C, whether we used 100,000 or 20,000 generations with early stopping.

### v2

@@ I Introduction
SDG techniques with built-in privacy assurances have however been shown to leak private information from the training data [5, 13, 19]. One fundamental way of showing leakages is to conduct a membership inference attack (MIA) on the synthetic data, the goal being to glean which records were present in the hidden training data [17, 39]. Practically speaking, a successful MIA on cancer-related synthetic data or an AI model could allow an attacker to determine whether an individual’s data was included in the training dataset. This could indirectly reveal that the individual had cancer or aid the attacker in performing more advanced attacks to infer aspects of their medical history.
@@ II-A Synthetic data generation
Many types of SDG algorithms have been developed to generate data in different ways. One prominent class of SDG employs generative adversarial networks (GANs). Algorithms using this paradigm perform quite well when generating images (e.g. [37, 41]). However, they tend not to produce high quality tabular data [29, 35]. The other main approach to SDG is to instead measure statistical properties from the training data $D_{train}$ and preserve them in the synthetic data $D_{synth}$. Specifically, we focus on privacy-preserving “marginals-based” SDG algorithms, i.e. those that maintain marginal probability distributions to keep this likeness. [^5] It is worth briefly mentioning that similar statistics-preserving SDG algorithms rely indirectly on marginals, via copulas to capture the dependencies between features. But we don’t study copulas algorithms here. Furthermore, some algorithms might be classified as projection-based [18]. However, many of these still operate by measuring marginals [2, 40] and using projection to preserve them. When it comes to tabular data, marginals-based SDG techniques are dominant in achieving the highest utility for machine learning applications while maintaining privacy [29, 35]. Naturally, a higher level of scrutiny must accompany their acclaim. New research has shown that diffusion-based generation of tabular data can produce data with impressive fidelity to the training data [22, 44, 38, 8]. However, as of this writing, most of them are not differentially private, and those that are have not been compared against differentially private marginals-based SDG.
In this paper, we use our novel MAMA-MIA attack to investigate vulnerabilities of three state-of-the-art SDG algorithms that employ marginals in some way: MST [26], PrivBayes [45], Private-GSD [23] (and we include an investigation on RAP [2] in the appendix). What is notable about these algorithms, unlike several popular others, is that they each provide robust differential privacy (DP) [9] guarantees to protect the training data. So to maximize the inherently-conflicting aims of achieving high quality and strong privacy, these algorithms use marginals and apply DP, respectively. Descriptions of MST, PrivBayes, RAP, and Private-GSD are included in Section  III-C.
@@
$\hskip-8.0pt\hat{P}_{D}(v_{c}\mid\mathbf{v})=\frac{|\{d\,|\,d\in D\wedge V_{c}(d)=v_{c}\ \wedge\mathbf{V}(d)=\mathbf{v}\}|}{|\{d\,|\,d\in D\wedge\mathbf{V}(d)=\mathbf{v}\}|}$ (2)
@@
Differential privacy (DP) The well-known definition of DP [9] states that a randomized algorithm $\mathcal{A}$ with range $O$ is $\varepsilon$-DP if for any two adjacent datasets $D_{1}$ and $D_{2}$ (datasets that differ in one entry)
@@ II-B Membership inference
Threat model Each MIA makes assumptions about what an adversary has access to, known as the “threat model”. Some MIAs use a threat model where an adversary knows little. Others grant a generous amount of knowledge to an adversary––the motivation being to drive the development of stronger privacy protections against stronger opponents. Our attack follows the latter inspiration. In our threat model, we use the “auxiliary data assumption”, which admits knowledge of the population data to the adversary [17]. This is a generally accepted assumption [39, 1], though is argued against in [12].
@@
An even stronger threat model could grant “white-box” knowledge, where an adversary knows the internal randomness (seed) of the target generator, and its weights learned during training on the hidden data. But we do not use this liberal assumption in our attack. There also is a distinction between having black-box knowledge of the generator to create a synthetic dataset from one’s own training data, not necessarily knowing the generator’s implementation (a slightly weaker assumption, used in [34]), and knowing its implementation (a stronger assumption, and used in [1]). We assume the latter here, even though most MIAs do not take advantage of this added information.
@@ II-C Related MIAs
References [4] and [34] describe attacks that rely on exhaustive “shadow-modelling” to learn about the hidden data. The idea behind these is to simulate the SDG algorithm on many random samples of the population data, thereby generating many “shadow datasets”, $\hat{D}_{synth}$. By generating many shadow synthetic datasets, it may be possible to notice differences in those that were trained using the target (i.e. the example the attacker is interested in) versus those that were not, informing whether or not the target was a member of the actual hidden training data $D_{train}$. Both these attacks use some kind of access to the model – either as prior knowledge of the SDG model in [34] or as query access to the model in [4] – to conduct the shadow modelling using one’s own training data. And both works show superior inference accuracy using this method. However, it tends to require generating several hundred shadow datasets, using scads of compute resources.
@@
The TAPAS library [^6] https://tapas-privacy.readthedocs.io/en/latest/index.html provides a framework for several types of MIAs based on different threat models [17], including shadow modelling such as “Groundhog” attacks [34], distance-based attacks [24, 42], probability estimation attacks [20, 15], and others. In both their analysis and in [27], shadow modelling attacks using random queries yield the highest accuracy when using the auxiliary data and black-box knowledge of the generator assumptions to generate shadow datasets. So in this paper, we provide a comparison of our attack with this leading MIA. (Meeus et al. find even greater success by smartly choosing which target records are most susceptible to attack [27]. We do not compare our work against this specialized approach, since ours is less restricted, and randomly samples which targets to attack.)
We also conduct a comparison of our attack against the DOMIAS MIA [39]. DOMIAS is highly relevant because, in addition to making clever use of the auxiliary data assumption (expounded upon in Section III), it outperforms several other MIAs [5, 13, 15]. But it does not use the black-box knowledge assumption, and is therefore generic to SDG algorithms.
A more detailed description of the MIAs we compare with in this work are given in Appendix A-F.
@@ III MAMA-MIA: the MArginals Measurment Aggregation-based Membership Inference Attack
MAMA-MIA constructs a novel density estimation $\zeta$ in place of $\Lambda$. $\zeta$ incorporates the normalization idea from DOMIAS, but is tailored to which marginals-based algorithm was used, and therefore has a much stronger ability to detect characteristics of the training data. Applicable to any marginals-based SDG algorithm, the heuristic behind MAMA-MIA is to:
@@ III-C Adaptation of MAMA-MIA to each SDG algorithm
Like MST, it uses the conditionals-based graphical approximation in order to generate $D_{synth}$, making this approximation highly relevant to an adversary. Edges that yield high mutual information are preferred by PrivBayes, but how many parents are allowed in each conditional depends on $\varepsilon$.
@@
A note on generalizability and efficiency Our density estimation approach can be tailored to similar SDG algorithms that estimate the probability distribution of a training dataset via marginals or other similar statistics, like conditionals or copulas. In Appendix section III-C, for example, we explain how MAMA-MIA can be adapted to RAP, another recent marginals-based SDG algorithm, and provide experimental results for it. Additionally, MAMA-MIA is highly efficient, and operates in linear time with respect to the size of $D_{synth}$ and $D_{aux}$ (we offer an asymptotic analysis in Appendix A-D). Not only does it not require generating hundreds of shadow datasets, but MAMA-MIA also has the benefit of operating on many targets at once without the runtime being noticeably impacted. That is, it is not necessary to train a new classifier per target.
@@ IV Experiments and results
We perform a series of experiments on MAMA-MIA to show its leading accuracy, and to demonstrate its far superior efficiency compared to state-of-the-art MIAs – TAPAS [17] and DOMIAS [39]. These experiments include attacking each SDG algorithm – MST, PrivBayes, and Private-GSD – under various configurations.
@@
A note on our TAPAS setup The TAPAS attack is set up to generate shadow datasets for one target at a time. However, in our experiments, we attack a set of targets, as is done in [1]. So we modify the TAPAS library to make shadow modelling much more efficient. We do this by making sure that all targets are in a different 50% of the shadow datasets, which allows us to reuse the shadow datasets in the attack for each target, thereby only requiring us to shadow model once, similar to the setup of [4].
While in [17] random queries are fed into the random forest classifier at the heart of the TAPAS attack, we feed it all possible 3-way queries. [^9] An “$n$-way query” is simply the count of rows in a dataset that match the target record’s values over a set of $n$ columns. In our investigation, using all queries either improved or did not affect the accuracy of TAPAS. This makes sense because a random forest will choose the best queries to use. Additionally, we explored using 1-way, 2-way, and 5-way queries. 3-way queries always yielded the highest accuracy.
Unless otherwise stated, we conduct all TAPAS attacks with 500 shadow datasets each.
@@ IV-A Experiment A results
Results are shown in Figure 3. The shaded regions depict +/- one standard deviation from the mean AUC. Our MAMA-MIA attack’s accuracy is at par with or better than the other MIAs on data generated by all three SDG algorithms. Not surprisingly, the accuracy improves when the privacy budget $\epsilon$ is larger, i.e. when the privacy guarantees offered by the DP SDG algorithms are lower. DOMIAS did not perform significantly above random guessing, neither with BNAF nor with KDE, the two density estimators used by Van Breugel at al. [39]. [^12] There were settings in which DOMIAS performed better than random guessing, but still far worse than MAMA-MIA, and only when the datasets used were small (see Appendix Section A-B). We also provide a tabled version of our full results, with standard deviations, as well as TPR / FPR curves when $\varepsilon=10$ in the Appendix A-E.
But the striking advantage our MAMA-MIA technique has over its fellow MIAs is its efficiency. Shown in Table I are the average runtimes of each of the MIAs in experiment A. These times include the full attack pipeline on the set of targets; that is, both the shadow modelling stage and the inference stage of the MAMA-MIA attack, and the TAPAS attack. Recall from Section III-A, we only shadow model 50 times for MAMA-MIA, where we terminate the SDG process early (as soon as the focal-points are determined). This is particularly relevant for SDG algorithms that are computationally intensive, like Private-GSD, where the TAPAS shadow modelling step entails 500 complete executions. [^13] The times given are processor time, and so can be shortened with parallel processing. For DOMIAS, there is no shadow modelling. But even so, when using BNAF, DOMIAS takes the longest (except for on Private-GSD).
TABLE I: Total pipeline runtimes for attacking all targets in one trial in experiment A, where $|D_{train}|=1,000$ records of SNAKE data.
@@ IV-D Additional results
We only conduct the TAPAS attack on Private-GSD for when $\varepsilon=10$ because of the high cost of generating 500 shadow datasets for each setting. To give a sense, just generating the Private-GSD shadow datasets for one TAPAS attack took 387 days of processor time. The second part of the attack that processes the shadow datasets took 19 minutes.
@@
The non-monotonicity of the qualities of Private-GSD housing data in Figure 9 part (C) is somewhat puzzling, and could be the result of an incompatibility between what our distance measurement captures and what Private-GSD captures on the binned housing data. The far-off distances in Private-GSD’s data in part (b) of the figure, and the odd spike in its distance in part (d), show how the quality of Private-GSD data degrades when only using 20,000 generations (the default amount) when $|D_{train}|$ and $|D_{synth}|\geq 3,000$ records. However, when 100,000 generations are used, the quality is as good as data produced by PrivBayes. We always allow for early stopping of Private-GSD, which can explain why, when $|D_{train}|$ and $|D_{synth}|\leq 1,000$ records, using 20,000 generations versus 100,000 generations yields data with the same distances.
So because of this, in all experiments except for experiment C, we set Private-GSD to use 100,000 generations.
$\textit{MA}=(\textit{tpr}-\textit{fpr}+1)/2$ (4)
where tpr, fpr are the true positive rate and false positive rate, computed by weighting individual predictions by their distance from a $0.5$ threshold: $2\cdot|0.5-p_{i}|,p_{i}\in P$. Like with AUC, an MA score of $0.5$ is as good as random guessing. In all of our experiments, MA almost always ended up being within $\pm$ 0.02 of what the AUC score was. Our accuracies are noticably better here than for the above experiments because the objective of this challenge was “set MI” (where targets are grouped into sets of five or more records, and either the entire set of records is in $D_{member}$ or none are). We provide more set MI results in Appendix A-E.
@@ V Discussion and broader impact
Acknowledgments
Steven Golob and Sikha Pentyala are Carwein-Andrews Distinguished Fellows. Steven Golob is supported by an NSF CSGrad4US fellowship. Sikha Pentyala is supported by a JP Morgan Chase PhD fellowship. This research was, in part, funded by the UW Global Innovation Fund, the UWT Founders Endowment and the National Institutes of Health (NIH) Agreement No.1OT2OD032581.
We would like to sincerely thank the organizers of the SNAKE challenge for posing this very interesting problem. Our research group, which has traditionally focused on privacy defenses, learned greatly from attempting to audit them.
@@ A-B MAMA-MIA on RAP
TABLE V: Runtimes of MAMA-MIA on SNAKE data when $\varepsilon=10$ (i.e. experiment B), separated into the shadow modelling stage, where we select FPs, and the inference stage, where we use them to compute $\zeta$. The runtime of the select FP stage is for 50 shadow modelling runs. The runtime of the compute $\zeta$ stage given is the average runtime for one trial, attacking all targets in $D_{target}$. Results for Private-GSD on 31,623 records are omitted because of computational limitations.
MAMA-MIA on MST MAMA-MIA on PrivBayes MAMA-MIA on Private-GSD
100 13.2 min 2.57 sec 2.11 sec 6.73 sec 2.57 min 16.3 sec
316 13.3 min 3.85 sec 2.63 sec 5.33 sec 2.69 min 38.8 sec
1,000 14.0 min 4.14 sec 6.46 sec 4.74 sec 4.93 min 79.9 sec
3,162 14.3 min 3.75 sec 15.7 sec 5.22 sec 8.09 min 2.17 min
10,000 14.5 min 3.02 sec 50.0 sec 4.37 sec 10.2 min 2.52 min
31,623 41.7 min 1.01 sec 4.37 min 4.43 sec N/a N/a
@@ A-C Experimental parameters
TABLE VI: Average AUC scores ($\pm$ one standard deviation) over 30 trials for MAMA-MIA and TAPAS attacks on the SNAKE data. $D_{train}=1,000$, and TAPAS uses 500 shadow datasets.
@@
TABLE VII: Average AUC scores ($\pm$ one standard deviation) over 30 trials for MAMA-MIA and TAPAS attacks on the housing data. $D_{train}=1,000$, and TAPAS uses 500 shadow datasets.
@@
For MST and PrivBayes, we use the implementations given in the Reprosyn library. [^17] https://github.com/alan-turing-institute/reprosyn. For Private-GSD and RAP, we use the authors’ implementations. [^18] https://github.com/giusevtr/private_gsd, [^19] https://github.com/amazon-science/relaxed-adaptive-projection We use the existing implementations of these DP SDGs. The main idea is to audit the existing algorithms and implementations [^20] We are aware of the bugs and incorrect implementations such as in PrivBayes as mentioned in [34]., and the assumptions that the meta data of the training set is available to the adversary.
For all experiments (except in experiment C because of computational limitations), we increase the default number of generations Private-GSD uses from 20,000 with early stopping to 100,000 with early stopping. As shown in Figure 9, this only makes a difference on the data quality for when $D_{train}$ and $D_{synth}$ have more than 1,000 rows. Therefore, this parameter change did not impact the accuracy of MAMA-MIA in experiment C, whether we used 100,000 or 20,000 generations with early stopping.